# Fenchel-Young Variational Inference

### v2

@@ 3.1 Fenchel-Young Variational Posteriors
Our proposed FYVI formulation is built by substituting the Kullback-Leibler divergence in (5) with the FY loss induced by a l.s.c. regularizer $\Omega$ with respect to a log-prior $\eta=\log p_{Z}$. This yields a new class of generalized posterior distributions of the following form:
$q_{x}^{\star}:=\arg\min_{q\in\mathcal{Q}}\underset{Z\sim q}{\mathbb{E}}\left[\ell(x;Z)\right]+L_{\Omega}(\eta;q)=\arg\min_{q\in\mathcal{Q}}\underset{Z\sim q}{\mathbb{E}}\left[\ell(x;Z)-\eta(Z)\right]+\Omega(q).$ (11)
At this point, we do not restrict or specify $\mathcal{Q}$.
@@
In this paper, we focus on FYVI with Tsallis $\rho$-negentropies $\Omega_{\rho}$ as regularizers, which as mentioned above, for $\rho>1$, naturally induce variational posteriors with adaptively sparse support, i.e., such that the support may be strictly smaller than $\mathcal{Z}$ and may depend on $x$. In contrast, for $\rho<1$, the variational posteriors may be denser (closer to uniform) than those obtained with standard Shannon entropy regularization (i.e., $\rho=1$); we will not consider this case, and leave it for future work.
@@
By definition, an exponential family has a support that cannot depend on its parameters [5]. Consequently, due to their variable support, generalized posteriors obtained via (11), where $\Omega=\Omega_{\rho}$, a Tsallis $\rho$-negentropy with $\rho>1$, cannot belong to an exponential family. This fact implies that contrary to the suggestion of the RoT (rule of three [34]), the variational family $\mathcal{Q}$ and the divergence cannot always be chosen independently of each other. Namely, adopting a divergence that may yield generalized posteriors with variable/adaptive support is incompatible with a variational family in any exponential family. In Appendix A, following similar derivations as Martins et al. [39, Prop. 8], we prove Proposition 1, demonstrating that for $\Omega=\Omega_{\rho}$, problem (11) is equivalent to a maximum Tsallis maximum $\rho$-entropy problem, the solution of which belongs to a $(2-\rho)$-exponential family,
$q_{x}^{\star}(z)$ $=\left[1+(1-\rho)\left(\eta(z)-\ell(x;z)-A_{\rho}\right)\right]_{+}^{\frac{1}{\rho-1}}=\exp_{2-\rho}\left[\eta(z)-\ell(x;z)-A_{\rho}\right],$ (12)
@@
where $A_{\rho}$ is a normalization function and $a]_{+}=\max\{0,a\}$ is the positive part operator, a.k.a. rectified linear unit (ReLU). Notice that the ReLU operation may produce zero results, thus assigning exactly zero probability (or probability density) to some elements of $\mathcal{Z}$, and the support of $q_{x}^{\star}$ may depend on $x$, something that is impossible in any exponential family.
@@
We stress that FYVI, with Tsallis $\rho$-negentropy regularization ($\rho>1$), leads to a qualitatively different type of sparsity from what is obtained by using sparsity-inducing priors or regularizers, such as LASSO—equivalently, Laplacian priors [52, 27]—or spike-and-slab priors [30]. Sparsity-inducing priors are used to encourage the estimated object—typically a vector of parameters of coefficients—to be sparse, but the corresponding posterior distributions have full support. For example, the Bayes posterior resulting from a Laplacian sparsity-inducing prior and a Gaussian observation model has full support. The reason why it is said to induce sparsity is that the posterior mode (the maximum a posteriori (MAP) estimate) can be exactly zero. In contrast, even using a non-sparsity-inducing Gaussian prior and a Gaussian observation model, together with FYVI with, say, Tsallis $2$-negentropy regularization, yields a generalized posterior density with bounded support, since it results from the application of the 2-entmax (sparsemax) prediction map [39].
@@ 3.2 Fenchel-Young Free Energy, Evidence, Evidence Lower Bound, and Posterior
By analogy with (4) the objective in (11) can be seen as a Fenchel-Young variational free energy (FYVFE). It can be useful to include an additional parameter $\beta$ weighting the KLD regularizer in (5) or the FY loss below, as in the $\beta$-variational auto-encoders ($\beta$-VAE [28]):
$F_{\ell,\Omega,\beta}(q,\eta;x)=\underset{Z\sim q}{\mathbb{E}}\left[\ell(x;Z)\right]+\beta L_{\Omega}(\eta;q).$ (13)
The parameter $\beta$ controls the strength with which the agreement between $q$ and $\eta$ is encouraged, and can also be seen from an information-theoretic perspective as controlling a rate-distortion trade-off [2]. In the sequel, omitting $\beta$ means assuming $\beta=1$.
By analogy with (3), the minimum of 13 is the (negative) FY evidence,
@@
and the minimizer $q_{x}^{\star}$ is the FY posterior. Since, by construction, the FY evidence $J_{\ell,\Omega}(x,\eta)$ is the maximum achieved by $F_{\ell,\Omega}(q,\eta;x)$, for any $q\in\Delta(\mathcal{Z}),$ the negative free energy $F_{\ell,\Omega}(q,\eta;x)$ is a FY evidence lower bound (FYELBO). For $\mathcal{Q}=\Delta(\mathcal{Z})$, $\Omega(q)=\Omega_{1}(q)=-H(q)$, and $\ell(x;z)=-\log p_{X|Z}(x|z)$, the FY evidence becomes the regular log-evidence $\log p_{X}(x)$ and the FY posterior recovers the standard Bayesian posterior $q_{x}^{\star}=p_{Z|X}(\cdot|x)$.
Although the ELBO (standard or FY) designation suggests that VI aims at bounding the evidence, the main goal of VI (standard or FY) is not to bound or approximate the evidence but rather to find the best possible approximation to the posterior in some variational family. In classical Bayesian analysis, a prior on the model space allows us to use the evidence for model comparison or hypothesis testing [37]. However, the FY evidence lacks the standard statistical interpretation as a marginal likelihood, and Bayesian model selection, validation, and hypothesis testing are not addressed here.
@@ 3.3 Fenchel-Young Expectation Maximization (FYEM)
Suppose $x$ contains $N$ independent observations, $x=(x_{1},...,x_{N})\in\mathcal{X}^{N}$, associated with a collection of latent variables $z=(z_{1},...,z_{N})\in\mathcal{Z}^{N}$, with corresponding variational distributions $q=(q_{1},...,q_{N})\in\Delta(\mathcal{Z})^{N}$. Let $\ell_{\theta}(x;z)=\sum_{i=1}^{N}\mathcal{L}_{\theta}(x_{i};z_{i})$ be a parameterized loss function, which is the sum of sample-wise losses, where $\theta$ is a vector of parameters to be estimated. The FYVI problem is to minimize the FYVFE (13) with respect to $q$ and $\theta$,
$\hat{q},\hat{\theta},\hat{\eta}$ $=\arg\min_{q,\theta,\eta}\sum_{i=1}^{N}F_{\ell_{\theta},\Omega}(q_{i},\eta;x_{i})=\arg\min_{q,\theta,\eta}\sum_{i=1}^{N}\Bigl{(}\underset{Z_{i}\sim q_{i}}{\mathbb{E}}\left[\mathcal{L}_{\theta}(x_{i};Z_{i})\right]+L_{\Omega}(\eta;q_{i})\Bigr{)}.$ (14)
where $\eta$ is a score function (possibly a log-prior), also to be estimated.
Alternating minimization w.r.t. $q$ and $(\theta,\eta)$ yields an EM-style algorithm: the E-step computes the posteriors of the latent variables; as in standard EM, each $q_{i}$ can be obtained independently of the others. The M-step updates the estimates of $\theta$ and $\eta$.
@@
$\hat{q}_{i}^{\mbox{\scriptsize new}}$ $=\arg\min_{q_{i}}\underset{Z_{i}\sim q_{i}}{\mathbb{E}}\left[\mathcal{L}_{\hat{\theta}}(x_{i};Z_{i})\right]+L_{\Omega}(\hat{\eta};q_{i})=\arg\max_{q_{i}}\underset{Z_{i}\sim q_{i}}{\mathbb{E}}\left[\hat{\eta}(Z_{i})-\mathcal{L}_{\hat{\theta}}(x_{i};Z_{i})\right]-\Omega(q_{i})$
$=\arg\min_{q_{i}}L_{\Omega}(\hat{\eta}-\mathcal{L}_{\hat{\theta}}(x_{i};\cdot);q)=\Pi_{\Omega}(\hat{\eta}-\mathcal{L}_{\hat{\theta}}(x_{i};\cdot)),$ (15)
for i=1,,Ni=1,...,N, where $\Pi_{\Omega}$ is the $\Omega$-regularized prediction map (7), namely softmax, sparsemax, or another $\rho$-entmax transformation.
@@
The M-step corresponds to updating $\hat{\theta}$ and $\hat{\eta}$, given $\hat{q}_{i}^{\mbox{\scriptsize new}}$. To update $\hat{\theta}$, we solve
$\hat{\theta}^{\mbox{\scriptsize new}}=\arg\min_{\theta}\sum_{i=1}^{N}\underset{Z_{i}\sim\hat{q}_{i}^{\mbox{\scriptsize new}}}{\mathbb{E}}[\mathcal{L}_{\hat{\theta}}(x_{i};Z_{i})],$
which depends on the particular form of $\mathcal{L}_{\hat{\theta}}$. To update $\hat{\eta}$, recall (see (6)) that $L_{\Omega}(\eta;q)=\Omega^{*}(\eta)-\mathbb{E}_{Z\sim q}[\eta(Z)]+\Omega(q)$, thus
$\hat{\eta}^{\mbox{\scriptsize new}}$ $=\arg\min_{\eta}\sum_{i=1}^{N}L_{\Omega}\bigl{(}\eta;\hat{q}_{i}^{\mbox{\scriptsize new}}\bigr{)}=\arg\min_{\eta}\Omega^{*}(\eta)-\frac{1}{N}\sum_{i=1}^{N}\underset{Z_{i}\sim\hat{q}_{i}^{\mbox{\scriptsize new}}}{\mathbb{E}}[\eta(Z_{i})]$
@@
where $\bar{q}=(1/N)\sum_{i=1}^{N}\hat{q}_{i}^{\mbox{\scriptsize new}}$, since the average of expectations under a set of probability (mass or density) functions is the expectation under the average of these functions. Assuming $\mathcal{Z}=\{1,...,K\}$ is a discrete set, $\eta$ is simply a vector $\eta=(\eta_{1},...,\eta_{K})\in\mathbb{R}^{K}$ and $q_{i}=(q_{i1},...,q_{iK})\in\Delta(\mathcal{Z})$ is probability distribution. In this case, $\mathbb{E}_{Z\sim\bar{q}}[\eta(Z)]=\langle\bar{q},\eta\rangle=\sum_{k=1}^{K}\bar{q}_{k}\eta_{k}$, thus
$\hat{\eta}^{\mbox{\scriptsize new}}\in\arg\min_{\eta}\;\Omega^{*}(\eta)-\langle\bar{q},\eta\rangle,$ (17)
where we have written the update as an inclusion because, as shown by Martins et al. [39], the minimizer may not be unique and is achieved by any $\hat{\eta}$ satisfying $\Pi_{\Omega}(\hat{\eta})=\bar{q}$. If $\Omega$ is the Shannon negentropy, $\Pi_{\Omega}$ is softmax, and $\hat{\eta}^{\mbox{\scriptsize new}}=(\log\bar{q}_{1},...,\log\bar{q}_{K})+C$, where $C$ is an arbitrary constant, which, without loss of generality, we set to zero, since it does not affect the algorithm. If $\Omega$ is the Tsallis 2-negentropy $\Omega_{2}$, $\Pi_{\Omega}(\eta)$ is the Euclidean projection of $\eta$ onto the simplex, and any $\hat{\eta}$ of the form $\hat{\eta}^{\mbox{\scriptsize new}}=(\bar{q}_{1},...,\bar{q}_{K})+C$, where $C$ is again an arbitrary constant that we set to zero, is a solution (although there may exist other solutions).

### v3

@@ 3.1 Fenchel-Young Variational Posteriors
FYVI problems are built by substituting the Kullback-Leibler divergence in (5) with the FY discrepancy induced by a regularizer $\Omega$ with respect to a log-prior $\eta=\log p_{Z}$. This yields a new class of generalized posterior distributions of the following form:
$q_{x}^{\star}$ $=\arg\min_{q\in\mathcal{Q}}\,\underset{z\sim q}{\mathbb{E}}\left[\ell(x;z)\right]+D_{\Omega}(\eta;q)$ (15)
$=\arg\min_{q\in\mathcal{Q}}\,\underset{z\sim q}{\mathbb{E}}\left[\ell(x;z)-\eta(z)\right]+\Omega(q).$ (16)
We first consider the case where $\mathcal{Q}=\Delta(\mathcal{Z})$ is the space of all probability distributions.
@@
We focus on FYVI with Tsallis $\rho$-negentropies $\Omega_{\rho}$ as regularizers, which for $\rho>1$ naturally induce variational posteriors with adaptively sparse support, that is, such that the support may be strictly smaller than $\mathcal{Z}$ and depends on observations $x$. In contrast, when $\rho<1$ the variational posteriors would be denser (closer to uniform) than those obtained with standard Shannon entropy regularization ($\rho=1$); we do not consider this case, and leave it for future work. Figure 1 illustrates the appeal of sparse posteriors on a 1D logistic regression problem: Due to the compact support, the approximate posteriors with $\rho=2$ and $\rho=1.5$ are closer to the MAP solution while still capturing uncertainty around it. Our FYVI framework gives us a new tool to control how close we want to be to the MAP solution, which we expect to be helpful in deep learning.
@@
By definition, exponential families ($\rho=1$) have support that cannot depend on their parameters [Barndorff-Nielsen]. Consequently, due to their variable support, generalized posteriors obtained through (16) where $\Omega$ is a Tsallis $\rho$-negentropy with $\rho>1$, cannot belong to an exponential family. This fact implies that contrary to knoblauch2022optimization’s rule of three suggestion the variational family $\mathcal{Q}$ and the divergence should not naively be chosen independently of each other. For example, adopting a divergence that may yield generalized posteriors with adaptive support is incompatible with a variational family in any exponential family. In Appendix A, following similar derivations as Prop. 8 in martins2022sparse, we show that for $\Omega=\Omega_{\rho}$ the optimization in (16) is equivalent to a maximum Tsallis $\rho$-entropy problem. The solution to that problem belongs to a $(2-\rho)$-exponential family of the form
$=\exp_{2-\rho}\left[\eta(z)-\ell(x;z)-A_{\rho}\right],$ (18)
@@
where $A_{\rho}$ is a normalization function and $a]_{+}=\max\{0,a\}$ is the positive part operator or rectified linear unit (ReLU). Notice that the ReLU operation may assign exactly zero probability (density) to some elements of $\mathcal{Z}$, and the support of $q_{x}^{\star}$ may depend on $x$, something that is impossible in any exponential family.
Figure 1: 1D logistic regression with a $\mathcal{N}(0,1)$ prior with $(2-\rho)$-Gaussian posteriors. Sparse posteriors (larger $\rho$) are closer to the MAP solution, as their truncation avoids sampling any points with high loss. In contrast, heavy-tailed posteriors move farther away from the high loss regions.
@@
We stress that FYVI, with Tsallis $\rho$-negentropy regularization and $\rho>1$ leads to a qualitatively different type of sparsity from what is obtained by using sparsity-inducing priors or regularizers, such as LASSO (equivalently, Laplacian priors [LASSO, Hastie]), or spike-and-slab priors [spike]. Sparsity-inducing priors are used to encourage the estimated object—typically a vector of parameters of coefficients—to be sparse, but the associated posterior distributions have full support. For example, the Bayes posterior resulting from a Laplacian sparsity-inducing prior and a Gaussian observation model has full support. The reason why it is said that such priors induce sparsity is that the posterior mode (the maximum a posteriori or MAP estimate) can be exactly zero. In contrast, using any non-sparsity-inducing Gaussian prior and a Gaussian observation model in FYVI with, say, Tsallis $2$-negentropy regularization, yields a generalized posterior with bounded support.
@@ 3.2 Fenchel-Young Free Energy, Evidence, Evidence Lower Bound, and Posterior
By analogy with classical VI (4) the objective in (16) can be interpreted as a Fenchel-Young variational free energy (FYVFE). It is useful to include an additional parameter $\beta$ weighting the KLD regularizer in (5) and the FY discrepancy below, as in $\beta$-VAEs [Higgins2017betaVAELB]:
$F_{\ell,\Omega,\beta}(q,\eta;x)=\underset{z\sim q}{\mathbb{E}}\left[\ell(x;z)\right]+\beta D_{\Omega}(\eta;q).$ (19)
The parameter $\beta$ controls the strength with which the agreement between $q$ and $\eta$ is encouraged, and can also be seen from an information-theoretic perspective as controlling a rate-distortion trade-off [alemi2018information]. In the sequel, omitting $\beta$ means assuming $\beta=1$.
By analogy with (3), the minimum of 19 is the (negative) FY evidence,
@@
and the minimizer $q_{x}^{\star}$ is the FY posterior. By construction, the FY evidence, $J_{\ell,\Omega}(x,\eta)$, is the maximum achieved by $F_{\ell,\Omega}(q,\eta;x)$, for any $q\in\Delta(\mathcal{Z})$. Therefore the negative free energy $F_{\ell,\Omega}(q,\eta;x)$ is a FY evidence lower bound (FYELBO). For $\mathcal{Q}=\Delta(\mathcal{Z})$, $\Omega(q)=\Omega_{1}(q)=-H(q)$, and $\ell(x;z)=-\log p_{X|Z}(x|z)$, the FY evidence reduces to the regular log-evidence, $\log p_{X}(x)$, and the FY posterior recovers the standard Bayesian posterior, $q_{x}^{\star}=p_{Z|X}(\cdot|x)$.
In classical Bayesian analysis, a prior on the model space allows us to use the evidence for model comparison or hypothesis testing [Bayes_factors]. While the FY evidence may lack the usual statistical interpretations such as a marginal likelihood, Bayesian model selection, validation, and hypothesis testing, we still expect it to be useful in practical settings. For instance, it may work well in deep learning, where the usual marginal likelihood has been successfully applied despite a lack of Bayesian modeling [immer2021scalable].
@@ 3.3 Fenchel-Young Expectation Maximization (FYEM)
Here, we show for the example of the expectation maximization algorithm how our new FYVI framework leads to the natural generalization of existing algorithms to accommodate flexible posterior distributions. We call the resulting method FYEM. Let $x$ be an $N$-dimensional vector of observations $x=(x_{1},...,x_{N})\in\mathcal{X}^{N}$, and $z$ a vector of associated latent variables $z=(z_{1},...,z_{N})\in\mathcal{Z}^{N}$, with corresponding variational distributions $q=(q_{1},...,q_{N})\in\Delta(\mathcal{Z})^{N}$. Let $\ell_{\theta}(x;z)=\sum_{i=1}^{N}\mathcal{L}_{\theta}(x_{i};z_{i})$ be a parameterized loss function: the sum of sample-wise losses, where $\theta$ is a vector of parameters to be estimated. Here, we minimize the FYVFE (19) with respect to $q$ and $\theta$,
$\hat{q},\hat{\theta},\hat{\eta}=\arg\min_{q,\theta,\eta}\,\sum_{i=1}^{N}F_{\ell_{\theta},\Omega}(q_{i},\eta;x_{i})$ (20)
$=\arg\min_{q,\theta,\eta}\,\sum_{i=1}^{N}\Bigl(\,\underset{z_{i}\sim q_{i}}{\mathbb{E}}\left[\mathcal{L}_{\theta}(x_{i};z_{i})\right]+D_{\Omega}(\eta;q_{i})\Bigr),$ (21)
and also with respect to $\eta$, which is a score function.
Alternating minimization w.r.t. $q$ and $(\theta,\eta)$ yields an EM-style algorithm: In the E-step we compute the posteriors of the latent variables; as in standard EM, each $q_{i}$ can be obtained independently of the others. In the M-step we update the estimates of $\theta$ and $\eta$.
@@
$\hat{q}_{i}^{\mbox{new}}$ $=\arg\min_{q_{i}}\underset{z_{i}\sim q_{i}}{\mathbb{E}}\left[\mathcal{L}_{\hat{\theta}}(x_{i};z_{i})\right]+D_{\Omega}(\hat{\eta};q_{i})$
$=\arg\max_{q_{i}}\underset{z_{i}\sim q_{i}}{\mathbb{E}}\left[\hat{\eta}(z_{i})-\mathcal{L}_{\hat{\theta}}(x_{i};z_{i})\right]-\Omega(q_{i})$
$=\arg\min_{q_{i}}D_{\Omega}(\hat{\eta}-\mathcal{L}_{\hat{\theta}}(x_{i};\cdot);q)$
$=\Pi_{\Omega}(\hat{\eta}-\mathcal{L}_{\hat{\theta}}(x_{i};\cdot)),$ (22)
for i=1,,Ni=1,...,N, where $\Pi_{\Omega}$ is the $\Omega$-regularized prediction map (11).
Figure 2: We compare qualitatively the standard, sparse, and hard versions of EM for GMM-based clustering. Data points are colored by their true labels, with x indicate noise points. Colored ellipses denote the data generating covariances; they are the same across the three panels. Black ellipses are level curves of the estimated GMM components and they illustrate the effect of E-step sparsity on the fit. Our sparse EM algorithm combines the best of both worlds: It is robust to the outliers while still retaining soft-assignments as in the original EM.
@@
Given $\hat{q}_{i}^{\mbox{new}}$, update $\hat{\theta}$ and $\hat{\eta}$. To update $\hat{\theta}$, we solve
$\hat{\theta}^{\mbox{new}}=\arg\min_{\theta}\sum_{i=1}^{N}\underset{z_{i}\sim\hat{q}_{i}^{\mbox{new}}}{\mathbb{E}}[\mathcal{L}_{\hat{\theta}}(x_{i};z_{i})],$
which depends on the particular form of $\mathcal{L}_{\hat{\theta}}$. To update $\hat{\eta}$, we recall from Eq. 7 that $D_{\Omega}(\eta;q)=\Omega^{*}(\eta)-\mathbb{E}_{z\sim q}[\eta(z)]+\Omega(q)$, and thus
$\hat{\eta}^{\mbox{new}}$ $=\arg\min_{\eta}\,\sum_{i=1}^{N}D_{\Omega}\bigl(\eta;\hat{q}_{i}^{\mbox{new}}\bigr)$
$=\arg\min_{\eta}\,\Omega^{*}(\eta)-\frac{1}{N}\sum_{i=1}^{N}\underset{z_{i}\sim\hat{q}_{i}^{\mbox{new}}}{\mathbb{E}}[\eta(z_{i})]$
@@
where $\bar{q}=(1/N)\sum_{i=1}^{N}\hat{q}_{i}^{\mbox{new}}$ by linearity of the expectation.
Assuming $\mathcal{Z}=\{1,...,K\}$ is a discrete set, $\eta$ is simply a vector $\eta=(\eta_{1},...,\eta_{K})\in\mathbb{R}^{K}$ and $q_{i}=(q_{i_{1}},...,q_{i_{K}})\in\Delta(\mathcal{Z})$ is a probability distribution. In this case, $\mathbb{E}_{z\sim\bar{q}}[\eta(z)]=\langle\bar{q},\eta\rangle=\sum_{k=1}^{K}\bar{q}_{k}\eta_{k}$, thus
$\hat{\eta}^{\mbox{new}}\in\arg\min_{\eta}\Omega^{*}(\eta)-\langle\bar{q},\eta\rangle,$ (24)
where we have written the update as an inclusion because, as shown by martins2022sparse, the minimizer may not be unique and is achieved by any $\hat{\eta}$ satisfying $\Pi_{\Omega}(\hat{\eta})=\bar{q}$. If $\Omega$ is the Shannon negentropy, then $\Pi_{\Omega}$ will be the softmax transformation, and $\hat{\eta}^{\mbox{new}}=(\log\bar{q}_{1},...,\log\bar{q}_{K})+C$, where $C$ is an arbitrary constant, which without loss of generality we can set to zero since it does not affect the algorithm. Under this choice, the FYEM recovers a regular EM algorithm. If $\Omega$ is the Tsallis 2-negentropy $\Omega_{2}$, then $\Pi_{\Omega}(\eta)$ is the Euclidean projection of $\eta$ onto the simplex, and any $\hat{\eta}$ of the form $\hat{\eta}^{\mbox{new}}=(\bar{q}_{1},...,\bar{q}_{K})+C$, where $C$ is again an arbitrary constant that we set to zero, is a solution of (24) (although there may exist others). If $\Omega=0$, we show in Appendix B that the above FYEM recovers the classical hard EM algorithm.